Charges 2uc. 4u C and 6u C are placed at the three corners A, B and C

respectively of a square ABCD of side x metre. Find, what charge must be

placed at the fourth corner so that the total potential at the centre of the

square is zero.

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Vnet = Va + Vb + Vc + Vd
 
Vnet = 0Vd = –( Va + Vb + Vc)Kq/r = –( K2/r + 4K/r + 6K/r )
here....K = 1/(4π£).
&
r is distance between center and corner
q = –(2+4+6) = –12mC